OpenStudy (anonymous):

Varia is studying abroad in Europe. She is required pay $3,500 (in US dollars) per year to the university, however, she must pay in euros. How many euros can Varia expect to pay per month to the university? (Round to the nearest whole euro.) 1.3687 euros = 1 US dollar

OpenStudy (anonymous):

1.3687 * 3500 = 4790.45 which rounds to 4790

OpenStudy (anonymous):

thank you so much allofus!

OpenStudy (anonymous):

Aneesha travels at a rate of 50 miles per hour. Morris is traveling 3 feet per second less than Aneesha. Which is the best estimate of the speed Morris is traveling? 1 mile = 5,280 feet

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Just a sec im figuring it out

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

,

OpenStudy (anonymous):

50 * 5280 = 264000 feet 264000/60 min = 4400feet / 1 min 4400-3 = 4397 4397*60 = 263820 feet 263820/50 = 5276.4mph Is this multiple choice?